# Specific heat of Gd3+ and Eu2+-based magnetic compounds

### v3

@@ VII Summary
We present a unified framework for understanding the specific heat of 4f7 magnetic compounds (Gd3+ and Eu2+), characterized by a ground state with $J=7/2$ and $L=0$. Through extensive quantum Monte Carlo simulations, we demonstrate that the specific heat is primarily governed by two parameters: the effective number of neighbors $z$ and axial anisotropy $K$.
@@
The effective number of neighbors $z$ quantifies the strength of spatial and quantum fluctuations, with larger $z$ values corresponding to reduced fluctuations and specific heat curves closer to mean-field predictions. Crucially, we find that the specific heat is remarkably insensitive to the detailed lattice structure for a given $z$, enabling a universal description of diverse magnetic systems.
@@
We successfully fit the specific heat of six experimental compounds (four Gd-based and two Eu-based) using only $z$ and $K$ as free parameters, with excellent agreement between theory and experiment. Notably, axial anisotropy is essential for Eu compounds but can be neglected for Gd compounds, reflecting the larger ionic radius of Eu2+ and its greater susceptibility to crystal-field effects. This work provides a simple yet powerful tool for interpreting specific heat data in 4f7 systems, significantly outperforming mean-field approaches and enabling the extraction of (effective) fundamental magnetic interaction parameters from experimental measurements. While our two-parameter model works well for the compounds studied, it may require additional parameters for systems with strong frustration or complex magnetic order.
Experimental researchers studying new 4f7 compounds can now follow this straightforward procedure: (1) Measure the specific heat and determine $T_{N}$; (2) Compare the $C/T$ vs $T/T_{N}$ curve to our reference data in Ref. [31]; (3) Identify the $z$ value that provides the best match; (4) Use Equation 4 and known crystal structure to constrain possible exchange interaction patterns. This approach provides physically meaningful constraints on magnetic interactions. These results could also guide the development of more accurate theoretical models for compounds.
@@ Acknowledgements.
The authors are grateful to I. Čurlík, L. de Sousa Silva, J.G.S. Duque, P.G. Pagliuso, and M. Avila for allowing us access to the original experimental data. A.A.A. acknowledges financial support provided by PICT 2020A 03661 of the Agencia I+D+i, Argentina. D.J.G. is supported by PIP 11220200101796CO of CONICET, Argentina. Computational resources were provided by the HPC cluster of the Physics Department at Centro Atómico Bariloche (CNEA), which is part of Argentina’s National High-Performance Computing System (SNCAD).
VIII Appendix
@@
$\langle x^{2}\rangle$ $=$ $J_{1}^{2}z_{1}\left[1+(z_{1}-1)(p-q)^{2}\right]+J_{2}^{2}z_{2}\left[1+(z_{2}-1)(r-s)^{2}\right]$ (14)

### v4

@@ VIII Summary
We present a unified framework for understanding the specific heat of 4f7 magnetic compounds (Gd3+ and Eu2+), characterized by a ground state multiplet $^{8}S_{7/2}$. Through extensive quantum Monte Carlo simulations, we demonstrate that the specific heat is primarily governed by two parameters: the effective number of neighbors $z$ and axial anisotropy $K$.
@@
The effective number of neighbors $z$ quantifies the strength of spatial and quantum fluctuations, with larger $z$ values corresponding to reduced fluctuations and specific heat curves closer to mean-field predictions. Crucially, we find that the specific heat is remarkably insensitive to the detailed lattice structure for a given $z$ as long as the couplings do not strongly increase with distance. This enables a universal description of diverse magnetic systems.
@@
We successfully fit the specific heat of six experimental compounds (four Gd-based and two Eu-based) using only $z$ and $K$ as free parameters, with excellent agreement between theory and experiment. Another Gd compound has been studied before [33]. The present approach allowed us to extract physical information not accessible through mean field. Notably, axial anisotropy is essential for Eu compounds but can be neglected for Gd compounds, reflecting the larger ionic radius of Eu2+ and its greater susceptibility to crystal-field effects.
This work provides a simple yet powerful tool for interpreting specific heat data in 4f7 systems, significantly outperforming mean-field approaches and enabling the extraction of (effective) fundamental magnetic interaction parameters from experimental measurements. While our two-parameter model works well for most of the compounds studied, it may require additional parameters for systems with strong frustration, complex magnetic order, or with interactions strengths that increases with distance as in the case of GdPtBi.
Experimental researchers studying new 4f7 compounds can now follow this straightforward procedure: (1) Measure the specific heat and determine $T^{*}$, the critical Néel or Curie temperature; (2) Compare the $(C/T)*T^{*}$ vs $T/T^{*}$ curve to our reference data in Ref. [31]; (3) Identify the $z$ value that provides the best match; (4) Use Eq. (4) and known crystal structure to constrain possible exchange interaction patterns. This approach provides physically meaningful constraints on magnetic interactions. These results could also guide the development of more accurate theoretical models for compounds.
@@ Acknowledgements.
The authors are grateful to I. Čurlík, L. de Sousa Silva, J.G.S. Duque, P.G. Pagliuso, and M. Avila for allowing us access to the original experimental data. D.J.G. is supported by PIP 11220200101796CO of CONICET, Argentina. Computational resources were provided by the HPC cluster of the Physics Department at Centro Atómico Bariloche (CNEA) and Clementina XXI. Language editing was supported by the use of Qwen3[43], with final approval of all content by the authors.
IX Appendix
@@
$\langle x^{2}\rangle$ $=$ $J_{1}^{2}z_{1}\left[1+(z_{1}-1)(p-q)^{2}\right]$ (14)
$+J_{2}^{2}z_{2}\left[1+(z_{2}-1)(r-s)^{2}\right]$